# Estimating Fold Changes from Partially Observed Outcomes with Applications in Microbial Metagenomics

### v1

@@ 4 Inference
We now discuss evaluating evidence for a difference in $\mathbb{E}W_{j}$ across levels of a covariate $X_{k}$, holding other covariates fixed. We explore the simple but common setting of testing $\overset{\circ}{H}_{0}:\beta_{kj}=0$, but the approach can be generalized to test more general hypotheses of the form $\mathbf{A}\beta=\mathbf{C}$ for fixed matrices $\mathbf{A}\in\mathbb{R}^{h\times p}$ and $\mathbf{C}\in\mathbb{R}^{h\times J}$. We are not able to test $\overset{\circ}{H}_{0}$ directly, as $\beta$ is not fully identified. However, we are able to test $H_{0}^{(g)}:\beta_{kj}=g(\beta_{k})$, or equivalently
$H_{0}:\text{ }\log\mathbb{E}[$ $W_{j}|X_{k}=x_{k},X_{-k}=x_{-k}]-\log\mathbb{E}[W_{j}|X_{k}=x_{k}+1,X_{-k}=x_{-k}]$
$=g(\log\mathbb{E}[W|X_{k}=x_{k},X_{-k}=x_{-k}]-\log\mathbb{E}[W|X_{k}=x_{k}+1,X_{-k}=x_{-k}])$
@@
where $S_{H_{0}}$ is the score evaluated at the maximum likelihood estimate $\hat{\beta}_{0}$ under the null, $I_{H_{0}}$ is a consistent estimate of the information matrix under the null, $F_{H_{0}}$ is given by $\frac{\partial h}{\partial\beta^{T}}$ evaluated at $\hat{\beta}_{0}$ and $D_{H_{0}}=\sum_{i=1}^{n}s_{i}(\hat{\beta}_{0})s_{i}(\hat{\beta}_{0})^{T}$ where $s_{i}(\beta)=\tilde{X}_{i}^{T}(Y_{i}-Y_{i\cdot}p_{i}(\beta))$ is the contribution to the score of $Y_{i}$. In practice, we use a score statistic $T^{\prime}_{RS}$ equal to $\frac{n}{n-1}T_{RS}$, where the term in front of $T_{RS}$ is an adjustment due to Guo et al. (2005) that reduces the small-sample conservatism of the test. Under $H_{0}$, $T^{\prime}_{RS}$ is asymptotically $\chi^{2}$-distributed with one degree of freedom; empirically, we observe that this test provides error rate control in large samples and conservative inference in small to moderate samples (Section 5). We also consider a robust Wald approach to testing hypotheses of the form $H_{0}^{(g)}:h(\beta)=0$. The robust Wald statistic is given by
@@
where $F_{H_{A}}$, $I_{H_{A}}$, and $D_{H_{A}}$ are as above but evaluated under the full model. Since robust Wald testing does not require refitting a model under a null hypothesis, it offers a computational advantage over robust score testing. However, as explored in Section 5, robust Wald tests fail to adequately control Type 1 error in small samples.
@@ 5 Simulations
We conduct a small simulation study to examine the Type 1 error and power of our proposed tests. We simulate in two distributional settings: a correctly specified setting in which counts are independently Poisson-distributed, and a misspecified setting in which the mean model is correctly specified but counts follow independent zero-inflated negative binomial (ZINB) distributions. To obtain draws from a ZINB distribution, we first draw from a Negative Binomial distribution with mean $\mu$ given by our model and dispersion parameter $\phi=5$ (thus, their variance is $\mu+\mu^{2}/\phi$). We multiply this draw with an independent Bernoulli draw $\xi$ with $Pr(\xi=1)=0.4$ to obtain our final ZINB draw.
Figure 2: Q-Q plots comparing empirical quantiles (y-axis) of the robust score (dark blue) and robust Wald (light blue) p-values to theoretical quantiles (x-axis). A conservative test will produce a curve above the line $x=y$ for small p-values and an anti-conservative test will produce a curve below it. To aid in comparison of lower quantiles, both axes are presented on a square-root scale. 10,000 simulations were performed.
@@
Figures 2 and 3 summarize the results of our simulations under the null and alternatives, respectively, with tabular results in Supporting Information Section 4. As expected, the robust score test exhibits conservatism in simulations under the null, with greater conservatism at smaller sample sizes and with ZINB-distributed counts. By contrast, the robust Wald test is anti-conservative in the smallest samples, with empirical Type 1 error 0.307 in ZINB simulations and 0.167 in Poisson simulations with $n=10$, $J=250$. In these two simulation settings, robust score tests have empirical Type 1 error 0.025 and 0.048, respectively. Notably, while the robust Wald test tends toward greater anti-conservatism when counts are not Poisson distributed, the robust score test appears to tend toward greater conservatism. At the highest sample size we considered, $n=250$, empirical Type 1 error is close to nominal regardless of distribution, number of taxa, or test.
Figure 3: Empirical power to reject $H_{0}$ at the 5% level, estimated from 500 simulations. The panels at left show power to detect a smaller effect ($\approx$ 3-fold ratio of mean concentrations), and the panels at right show power to detect a larger effect ($\approx$ 150-fold ratio of mean concentrations). See Supporting Information Figure 1 for power curves.
@@
In simulations under the alternative, we observe higher power when counts are drawn from a Poisson than from a ZINB distribution, and when sample size is larger. While the robust Wald test exhibits higher power than the robust score test at smaller sample sizes, it also fails to control Type 1 error, so we nonetheless do not recommend this test for use in small or moderate samples. We also observe a modest impact of total number of taxa $J$ on power: for ZINB simulations, we observe nondecreasing power in $J$. However, power of the score test does appear to decrease in $J$ for Poisson simulations with $n=10$.
@@ 6 Data analysis: Colorectal cancer microbial metaanalysis
In this section we consider a shotgun metagenomic dataset published by Wirbel et al. (2019). This dataset comprises a meta-analysis of five studies, each targeting possible associations between the composition of the human gut microbiome and colorectal cancer (CRC) risk. In each study, fecal samples were taken from participants with colorectal cancer as well as from healthy controls. We are interested in identifying microbial strains unusually enriched or depleted in study participants with colorectal cancer as compared to otherwise similar control participants.
@@
Each observation $Y_{ij}\in\mathbb{R}_{\geq 0}$ reflects the observed abundance of microbial strain $j$ in fecal material of study participant $i$, measured as the depth of coverage of marker genes calculated by the mOTU profiler v2 (Milanese et al., 2019). We consider the “ideal” but unobserved data $W_{ij}$ to be the cell concentration of strains in the stool of study participants. Participant data was pooled from five studies, yielding 575 study participants among whom 2342 unique strains were detected. In addition to metagenomic data, Wirbel et al. (2019) published clinical and demographic metadata including colorectal cancer status, age in years, binary gender, BMI, and timing of sampling relative to colonoscopy (sample taken prior to or following colonoscopy). Wirbel et al. (2019) applied stringent filtering criteria to remove low-abundance strains from their dataset. While not necessary for our methodology, to enable the comparison of our results with theirs, we applied the same filtering criterion, excluding all mOTUs that failed to exceed empirical relative abundances of $10^{-3}$ in at least three studies. We excluded 9 participants with missing covariate data (8 for BMI and 1 for timing of sampling relative to colonoscopy). With these exclusions, we analyze $n=566$ study participants and $J=845$ metagenomic strains.

### v2

@@ 5 Inference
We now discuss evaluating evidence for a difference in $\mathbb{E}W_{\cdot j}$ across levels of a covariate $X_{\cdot k}$, holding other covariates fixed. We explore the simple but common setting of testing $\overset{\circ}{H}_{0}:\beta_{kj}=0$, but the approach can be generalized to test more general hypotheses of the form $\mathbf{A}\beta=\mathbf{C}$ for fixed matrices $\mathbf{A}\in\mathbb{R}^{h\times p}$ and $\mathbf{C}\in\mathbb{R}^{h\times J}$. We are not able to test $\overset{\circ}{H}_{0}$ directly, as $\beta$ is not fully identified. However, we are able to test $H_{0}^{(g)}:\beta_{kj}=g(\beta_{k})$, or equivalently
$H_{0}:\text{ }\log\mathbb{E}[W_{\cdot j}|X_{\cdot}$ $=\mathbf{x}_{x_{k}+1}]-\log\mathbb{E}[W_{\cdot j}|X_{\cdot}=\mathbf{x}]$
$=g\left(\log\mathbb{E}[W_{\cdot j}|X_{\cdot}=\mathbf{x}_{x_{k}+1}]-\log\mathbb{E}[W_{\cdot j}|X_{\cdot}=\mathbf{x}]\right)$
@@
where $S_{H_{0}}$ is the score evaluated at the maximum likelihood estimate $\hat{\beta}_{H_{0}}$ under the null, $I_{H_{0}}$ is a consistent estimate of the information matrix under the null, $F_{H_{0}}$ is given by $\frac{\partial h}{\partial\beta^{T}}$ evaluated at $\hat{\beta}_{H_{0}}$ and $D_{H_{0}}=\sum_{i=1}^{n}s_{i}(\hat{\beta}_{H_{0}})s_{i}(\hat{\beta}_{H_{0}})^{T}$ where $s_{i}(\beta)=\tilde{X}_{i}^{T}(Y_{i}-\big{|}\big{|}Y_{i}\big{|}\big{|}_{1}p_{i}(\beta))$ is the contribution to the score of the $i$th observation. In practice, we use a score statistic $T^{\prime}_{RS}$ equal to $\frac{n}{n-1}T_{RS}$, where the term in front of $T_{RS}$ is an adjustment due to Guo et al. (2005) that reduces the small-sample conservatism of the test. Under $H_{0}$, $T^{\prime}_{RS}$ is asymptotically $\chi^{2}$-distributed with one degree of freedom; empirically, we observe that this test provides error rate control in large samples and conservative inference in small to moderate samples (Section 6). We also consider a robust Wald approach to testing hypotheses of the form $H_{0}^{(g)}:h(\beta)=0$. The robust Wald statistic is given by
@@
where $F_{H_{A}}$, $I_{H_{A}}$, and $D_{H_{A}}$ are as above but evaluated under the full model. Since robust Wald testing does not require refitting a model under a null hypothesis, it offers a computational advantage over robust score testing. However, as explored in Section 6, robust Wald tests fail to adequately control Type 1 error in small samples. In contrast, robust score tests controlled error rates in all investigated scenarios.
Figure 2: Q-Q plots comparing empirical quantiles (y-axis) of the robust score (dark blue) and robust Wald (light blue) p-values to theoretical quantiles (x-axis). P-values from ALDEx2 (orange), ANCOM-BC2 (pink), a t-test on centered log ratio (CLR) transformed abundances (brown), and DESEq2 (green) are also compared. A conservative test will produce a curve above the line $x=y$ for small p-values and an anti-conservative test will produce a curve below it. To aid in comparison of lower quantiles, both axes are presented on a square-root scale. 500 simulations were performed.
Figure 3: Empirical power to reject $H_{0}$ at the 5% level (y-axis) across a range of effect sizes (x-axis), estimated from 500 simulations. $\beta_{1}=1$ represents an $\approx$ 3-fold difference in the ratio of true abundances across groups while $\beta_{1}=5$ represents an $\approx$ 150-fold ratio. Only valid tests (those that control Type 1 error rates for the given sample size, number of categories and data distributions) are shown.
@@ 6 Simulations
We conduct a simulation study to examine the Type 1 error and power of our proposed tests. We simulate in two distributional settings: a correctly specified setting in which counts are independently Poisson-distributed, and a misspecified setting in which the mean model is correctly specified but counts follow independent zero-inflated negative binomial (ZINB) distributions. To obtain draws from a ZINB distribution, we first draw from a Negative Binomial distribution with mean $\mu$ given by our model and dispersion parameter $\phi=5$ (thus, their variance is $\mu+\mu^{2}/\phi$). We multiply this draw with an independent Bernoulli draw $\xi$ with $Pr(\xi=1)=0.4$ to obtain our final ZINB draw.
@@
Figures 2 and 3 summarize the results of our simulations under the null and alternatives, respectively, with tabular results for the proposed methods in Supporting Information (Section 12). As expected, the robust score test exhibits conservatism in simulations under the null, with greater conservatism at smaller sample sizes and with ZINB-distributed counts. By contrast, the robust Wald test is anti-conservative in the smallest samples, with empirical Type 1 error $0.316$ in ZINB simulations and $0.160$ in Poisson simulations with $n=10$, $J=250$. In these two simulation settings, robust score tests have empirical Type 1 error $0.034$ and $0.044$, respectively. Notably, while the robust Wald test tends toward greater anti-conservatism when counts are not Poisson distributed, the robust score test appears to tend toward greater conservatism. At the highest sample size we considered, $n=250$, empirical Type 1 error is close to nominal regardless of distribution, number of taxa, or proposed test.
@@
In simulations under the alternative, we observe higher power when counts are drawn from a Poisson than from a ZINB distribution, and when the sample size is larger. While the robust Wald test exhibits higher power than the robust score test at smaller sample sizes, it also fails to control Type 1 error, so we do not recommend this test for use in small or moderate samples. We also observe a modest impact of total number of taxa $J$ on power: for ZINB simulations, we observe nondecreasing power in $J$. However, power of the score test does appear to decrease in $J$ for Poisson simulations with $n=10$.
We also compare error rates of our proposed methods to popular differential abundance methods that test comparable null hypotheses (Figures 2 and 3). Three of the most widely-used methods (ALDEx2 (Fernandes et al., 2014), ANCOM-BC2 (Lin and Peddada, 2024), and a two-sample t-test on CLR-transformed counts) add a small non-zero “pseudocount” $\epsilon$ to all values prior to log-transformations. In part because $\log\left(\mathbb{E}Y_{ij}\right)\neq\mathbb{E}\left[\log\left(Y_{ij}+\epsilon\right)\right]$, these methods do not test identical null hypotheses to the proposed method. However, the null hypothesis tested by these methods holds (for data generating mechanisms considered) when $\beta_{1}=0$, and does not hold when $\beta_{1}\neq 0$. Accordingly, we treat as comparable the nulls tested by the proposed method, ALDEx2, ANCOM-BC2, a two-sample t-test on CLR-transformed counts, as well as DESeq2 (Love et al., 2014). The null hypothesis does not necessarily coincide for corncob (Martin et al., 2020) or IFAA (Li et al., 2021), and therefore we do not consider these methods in this comparison.
In most settings investigated, ALDEx2 has conservative Type 1 error rates, resulting in its relatively low power compared to the proposed score test, though ALDEx2 can be anticonservative, with 5% error rates as high as 28% when $n=250$, $J=250$, and $Y$ follows a Poisson distribution. ANCOM-BC2 generally controls error rates when the observed counts follow a Poisson distribution, with the exception of $n=250$ and $J=250$, when its error rate at the 5% level is 32%. It also fails to control error rates in most ZINB-distributed count settings. Because methods that do not control error rates are not valid hypothesis tests ($\text{Pr}(p\leq\alpha)\leq\alpha$ under the null defines a hypothesis test), we only compare power amongst methods that control error rates. Accordinly, we find that ANCOM-BC2 has higher power than the proposed method for ZINB-distributed counts when $n=10$. A two-sample t-test on CLR-transformed counts (with a pseudocount of 1 added to all $Y_{ij}$ values) controls error rates in most ZINB settings, but not in Poisson settings. When error rates are controlled, the CLR t-test has lower power than the proposed method. DESeq2 typically fails to control the Type 1 error rates, but has high power in the few cases where errors are controlled (Poisson data with $n=10$). In contrast, the proposed score test controls error rates in all settings, and consistently displays high power. The power of the proposed score test can be outperformed (with respect to power) by the proposed Wald test and ANCOM-BC2, but these methods do not consistently control error rates. Overall, we find the proposed score test controls error rates across a wide variety of data generating processes, sample sizes, and species diversities; and high power to reject false null hypotheses when compared to rate-controlling alternatives.
@@ 7 Data analysis: Colorectal cancer microbial metaanalysis
We now investigate microbial strains that are unusually enriched or depleted in populations with colorectal cancer (CRC) compared to otherwise similar control populations. We consider a shotgun metagenomic dataset published by Wirbel et al. (2019), which is a meta-analysis of five studies (total $n=575$), each of which obtained fecal samples from participants with colorectal cancer and cancer-free controls. We consider the “ideal” but unobserved data $W_{ij}$ to be the cell concentration of microbial strain $j$ in fecal material of study participant $i$, and let $Y_{ij}\in\mathbb{R}_{\geq 0}$ be the observed abundance measured as the depth of coverage of marker genes calculated by the mOTU profiler v2 (Milanese et al., 2019). 2342 unique strains were detected. In addition to metagenomic data, Wirbel et al. (2019) published clinical and demographic metadata including colorectal cancer status, age in years, binary gender, BMI, and timing of sampling relative to colonoscopy (sample taken prior to or following colonoscopy). Wirbel et al. (2019) applied stringent filtering criteria to remove low-abundance strains from their dataset. While not necessary, to enable the comparison of our results with theirs, we applied the same filtering criterion (excluding all taxa that did not exceed empirical relative abundances of $10^{-3}$ in at least three studies). We excluded 9 participants with missing covariate data (8 for BMI and 1 for timing of sampling relative to colonoscopy). With these exclusions, we analyze $n=566$ study participants and $J=845$ metagenomic strains.
We fit our proposed model, applying Firth penalization as well as a smoothed median identifiability constraint to the rows of $\beta$. We include, in addition to an intercept and a term for CRC status (1 for CRC and 0 otherwise), terms in our model for age (linear spline with a single knot at median age 64 years), BMI (linear spline with a single knot at median BMI 24.7 kg/m2), gender (1 for Male and 0 otherwise), study, and whether fecal samples were taken after colonoscopy (1 if after, 0 if prior). Thus, coefficients on CRC status estimate the difference in log mean cell concentration comparing CRC patients to non-CRC controls who are alike in gender, BMI, age, study, and timing of stool sampling relative to colonoscopy, relative to typical such differences across taxa. We perform robust score tests and control the false discovery rate (using the Benjamini and Hochberg (1995) (BH) procedure, following Wirbel et al. (2019)) at $0.1$.
Robust Wald confidence intervals for the coefficients on CRC status are shown in Figure 4. We identify 30 taxa for which we reject the null hypothesis of equal log mean cell concentration for CRC patients and (otherwise alike) non-CRC controls at a false discovery rate of 0.1 based on the robust score test. These taxa are denoted by navy-colored confidence intervals. Taxa are ordered along the x-axis by the proportion of participants in which they are detected, highlighting that significantly differentially abundant taxa are identified across the spectrum of detection levels. Taxa our analysis identifies as more enriched than typical among CRC patients relative to similar controls include an unknown Dialister species estimated to have ratio of mean cell concentration 49 times greater than typical (95% CI $12-210$; $p=1.1\times 10^{-6}$), Porphyromonas uenonis (30 times greater; 95% CI $6.2-150$; $p=0.0025$), Fusobacterium nucleatum s. animalis (24 times greater; 95% CI $11-52$; $p=2.2\times 10^{-5}$), and Gemella morbillorum (23 times greater; 95% CI $7.6-69$; $p=0.0016$). Confidence intervals for all effect sizes that are significantly associated with CRC status at FDR cutoff $0.1$ are given in Figure 5.
We identify a number of taxa associated with CRC status that are detected in relatively high proportion in the study populations. For example, we estimate for B. kashiwanohense, which was detected in 35.7% of study participants, that the ratio of mean stool cell concentrations for this taxon comparing CRC patients to alike controls is 8.6 times lower (95% CI 3.6 – 20.9; unadjusted robust score $p=0.0022$) than the typical such ratio among taxa included in our analysis. This is consistent with a possible protective effect of B. kashiwanohense, but we cannot rule out association due to unmeasured confounding (e.g., through diet) or reverse causation (e.g., development of CRC induces relative depletion of B. kashiwanohense). We note that B. kashiwanohense was not identified as a taxon negatively associated with CRC status in the analysis of Wirbel et al. (2019).
Figure 6: Negative $\log_{10}$ p-values from distinct approaches to differential abundance analyses based on the Wirbel et al. (2019) dataset. The proposed method’s robust score test p-values are shown on the x-axis, and comparator methods on the y-axis. Wirbel et al. (2019) report some p-values equal to zero; to allow graphing on the log scale, we replace these values with $10^{-25}$. In addition to the missing/uncomputable p-values of corncob, DESeq2 and IFAA, two DESeq2 p-values and three ANCOM-BC2 p-values were so small as to make plotting infeasible; these observations correspond to $(p_{\text{radEmu}},p_{\text{DESeq2}})$ tuples of $(10^{-1.8},10^{-31})$ and $(10^{-2.8},10^{-33})$ and $(p_{\text{radEmu}},p_{\text{ANCOM-BC2}})$ tuples of $(10^{-1.8},10^{-26})$, $(10^{-3.7},10^{-39})$, and $(10^{-1.9},10^{-26})$.
We compare our results to those of Wirbel et al. (2019), as well as to 5 other popular differential abundance methods in the microbiome literature, in Figure 6. All panels show robust score p-values obtained from our adjusted model on the x-axis, and p-values obtained from comparator methods on the y-axis. We briefly describe details for each method here; see the Code Appendix for a reproducible workflow. (Top left) corncob (Martin et al., 2020) likelihood ratio test p-values with design matrix as for the proposed method and timing of sampling and study as dispersion covariates; (top center) ALDEx2 (Fernandes et al., 2014) with centered log-ratio transform and design matrix as for the proposed method and 512 Monte Carlo iterations; (top right) DESeq2 (Love et al., 2014) likelihood ratio test with design matrix as for the proposed method and “local” estimation of dispersions; (lower left) IFAA (Li et al., 2021) with design matrix as for the proposed method; (lower center) ANCOM-BC2 (Lin and Peddada, 2024) with design matrix as for the proposed method and without prevalence filtering; and (lower right) the results of Wirbel et al. (2019), who report a blocked Wilcoxon test performed on proportion-scale mOTU data, with blocking on study and timing of sampling relative to colonoscopy (five p-values of zero are altered to $10^{-25}$ for visual comparison). The proposed method produces smaller p-values in 44%, 68%, 52%, 49%, 20% and 41% of cases for corncob, ALDEx2, DESeq2, IFAA, ANCOM-BC2 and blocked Wilcoxon respectively. Thus, consistent with the results of our simulation study (ZINB, $J=250$ and $n=250$), our method typically produces smaller p-values than ALDEx2 (which was generally found to be conservative) but larger p-values than ANCOM-BC2 (which was generally found to be anti-conservative). In general, we observe weak correlations between all methods’ p-values, consistent with documented discrepancies between widely used differential abundance methods in microbiome science (Nearing et al., 2022). That said, our semiparametric estimation procedure, model-robust inference method, and simulation study results give confidence that our proposed method presents reliable inference with high power.
We also contrast our results with those of Wirbel et al. (2019), finding substantial variation in the ordering of p-values between our proposed method and the blocked Wilcoxon test (Spearman correlation $0.44$). Only 45 of the 94 taxa Wirbel et al. (2019) identified as significantly associated with CRC status are among the 94 taxa most significantly associated with CRC status in our analysis, and of the 30 taxa we highlight at FDR level $0.1$, only 10 are among the 30 most highly significant taxa identified by Wirbel et al. (2019). We believe that this lack of concordance is likely due to two major factors. Firstly, our model adjusts for a larger set of potential confounders than do Wirbel et al. (2019), which may account for our discrepant findings. In addition, we estimate effects on the scale of log mean cell concentration, whereas Wirbel et al. (2019) conduct tests on the read proportion scale. Because a perturbation to the abundance of a small number of taxa on the cell concentration scale induces a dense perturbation on the proportion scale, analyses on the proportion scale may highlight taxa that do not vary in a scientifically interesting way on the cell concentration scale. Similarly, analyses performed on proportion data may fail to detect scientifically meaningful effects on the cell concentration scale because of fluctuations in cell concentration among other taxa.
While it is beyond the scope of the present study to experimentally validate which taxa are truly differentially abundant in colorectal cancer gut microbiomes, we investigate the empirical number of false discoveries of all methods applied to this dataset using a permutation approach. We create 20 datasets $(\mathbf{X}^{*},\mathbf{Y}^{*})$ with no association between microbial abundances and cancer diagnosis by randomly reordering the rows of the design matrix $\mathbf{X}$ such that the permuted design matrices $\mathbf{X}^{*}$ satisfies $\text{Pr}(X_{i}^{*}=X_{i^{\prime}})=1/n$ for all $i,i^{\prime}$, but $\mathbf{Y}=\mathbf{Y}^{*}$. This approach leaves the joint distribution of the covariates unchanged but guarantees conditional independence between $Y_{i\cdot}^{*}$ and $X_{i,\text{CRC}}^{*}$, thereby guaranteeing that $\beta_{\text{CRC},j}^{*}=0$ for all $j$. We ran all analyses as before on the permuted 20 datasets, and report the distribution of the number of taxa that each method determines to be differentially abundant when controlling the false discovery rates at 10% (the empirical false discovery rate) in SI Table 3. Remarkably, our proposed method (radEmu score test) and ALDEx2 make no false discoveries across any of the 20 datasets. ANCOM-BC2 has the highest discovery proportion in this null dataset (median 37%; IQR: 36–39%), followed by the radEmu Wald test (which we do not recommend in general; median 3.2%; IQR: 2.6–3.6%), DESeq2 (median 0.5%; IQR: 0.2–1.1%), IFAA (median 0%; IQR: 0–0%; max 0.2%), and the CLR t-test (median 0%; IQR: 0–0%; max 0.1%). These results align with our simulation study, which demonstrated consistent error rate control of the proposed score test, very conservative behavior of ALDEx2, and anti-conservative behavior of the t-test, DESeq2 and ANCOM-BC2 for $n=250,J=250$ and zero-inflated Negative Binomially-distributed data. That said, we found DESeq2 to make fewer false discoveries than ANCOM-BC2, unlike in the the simulation study, which may be attributable to the differences between the ZINB distribution and the (unknown) true data generating process. ANCOM-BC2 also reports if discoveries pass a “sensitivity score filter”, but even with this sensitivity filter, a median of 14 discoveries were made (median 1.6%; IQR: 1.3–2.0%; max 3.3%). Combining these results with the findings of Section 6, we find DESeq2 and ANCOM-BC2 to be anticonservative in general (consistent with the benchmarking study of Cho et al. (2023)), and similarly for the CLR t-test but to a lesser degree. We find the performance of the proposed method to be most similar to that of ALDEx2, with the proposed method having higher power than ALDEx2 without sacrificing error rate control.
8 Discussion
In this paper we present a regression method for inference on means of a nonnegative multivariate outcome $W$ observed after perturbation by unknown sample-specific scaling terms $z$ and category-specific detection effects $\delta$. Our method is motivated by microbiome sequencing experiments, in which outcomes of interest (microbial cell concentration across many species) are typically observed indirectly via sequencing, which is known to be subject to both sample-specific scaling and taxon-specific distortion. However, we anticipate that our method may find application in other settings, especially in other ’omics fields where similar scaling and detection effects are likely to distort measurements.
Our method has a number of advantages over existing methods commonly applied to perform differential abundance analyses in microbiome studies. Firstly, it targets a population estimand that can be distinguished from sample- and category-scaling effects and that exists under realistic conditions (positivity of population mean abundances $\mathbb{E}W_{\cdot j}$). Notably, we do not require positivity of true abundances $W_{ij}$ nor observed abundances $Y_{ij}$. Secondly, our method does not require data transformations, such as log-ratio transformations or relative abundance transformations. This allows us to use information regarding the precision of observations in estimation. It also avoids ad-hoc data manipulation (e.g., adding a small pseudocount to zero observations, restricting attention to nonzero counts, or infering whether observed zeroes are due to insufficient sequencing depth) that is necessary to apply log-ratio transformations to data containing zeroes. Additionally, we implement a flexible inference approach that controls Type 1 error even in small samples with large excess-Poisson dispersion and high degrees of sparsity, and observed that the approach made no false discoveries on a shotgun sequencing dataset with true signal removed. In a simulation study, the proposed score test was the only method with error rate control in all data settings, and had comparable power to other error rate-controlling methods in most of these settings. Finally, we incorporate an identifiability constraint that may be flexibly specified to suit varying scientific settings. In particular, we do not require a reference taxon (though if available, the reference taxon can define the identifiability constraint; see Example 1) nor the assumption of a sparse signal (see Example 2).
Our method is not without limitations, and admits several possible extensions. While our target estimand is identifiable in the presence of systematic measurement error in the form of multiplicative row and column perturbations to its mean, no such guarantees exist for other forms of measurement error. A notable form of measurement error that is not addressed by our method is cross-sample contamination. Accordingly, we encourage the use of decontamination techniques (Davis et al., 2018; Clausen and Willis, 2022) prior to the use of our proposed method. In addition, in contexts where it is scientifically plausible that the parameter of interest $\beta_{k}$ is sparse, enforcing sparsity via $\ell_{1}$ penalization is an attractive option. However, we note that if most elements of $\beta_{k}$ are truly equal to zero, $\beta_{k}$ is correctly specified under pseudo-Huber constraints up to negligible differences between medians and pseudo-Huber centers (see Example 2). Furthermore, while estimation will be valid in the presence of longitudinal or cluster-dependent observations, it remains to extend our inferential methods to this setting. Finally, computation can become prohibitively expensive when $J\gtrapprox 10^{4},$ and exact and approximate algorithms could be explored to enhance the scalability of the method.
9 SI: Proofs of identifiability and asymptotic results
Here, we provide formal proofs of the results stated in the text.
Lemma 3.
For arbitrary $b\in\mathbb{R}^{p\times J}$, define
$G_{b}=\{\beta^{\star}\in\mathbbm{R}^{p\times J}:\beta^{\star}=b+\alpha\mathbf{1}_{J}^{T}\text{ for some }\alpha\in\mathbb{R}^{p}\}.$
If $\mathbf{X}$ has full column rank, for any $\beta^{\star},\beta^{\dagger}\in\mathbb{R}^{p\times J}$ there exist $\mathbf{z}\in\mathbb{R}^{n}$ and $\mathbf{z}^{\dagger}\in\mathbb{R}^{n}$ such that under model (3), $\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\star},\mathbf{z}]=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\dagger},\mathbf{z}^{\dagger}]$ if and only if $\beta^{\dagger}\in G_{\beta}$.
Proof.
Suppose $\beta\in G_{\beta}$. Then there exists an $\alpha\in\mathbb{R}^{p}$ such that $\beta=\beta^{\star}+\alpha\mathbf{1}_{J}^{T}$ and by ((4), main text), for any $\mathbf{z}$ and $\mathbf{z}^{\star}:=\mathbf{z}+\mathbf{X}\alpha$, $\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\star},\mathbf{z}]=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta,\mathbf{z}^{\star}]$. Now suppose $\beta\not\in G_{\beta^{\star}}$. If there exist $\mathbf{z},\mathbf{z}^{\star}$ such that $\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\star},\mathbf{z}]=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta,\mathbf{z}^{\star}]$, we have
$\mathbf{z}\mathbf{1}_{J}^{T}+\mathbf{X}\beta^{\star}=\mathbf{z}^{\star}\mathbf{1}_{J}^{T}+\mathbf{X}\beta$